Grade 12Physical Chemistry

at 10​​ degree celcius the osmotic pressure of urea solution is 500mm. the solution is diluted and the temperation is raised to 25 degree celcius ,when the osmotic pressure is found to be 105.3mm.the extend of dilution is?

shaily
9 Years agoGrade 12

1 Answer

saipallavi
9 Years ago
at 10​​ degree celcius the osmotic pressure of urea solution is 500mm. the solution is diluted and the temperation is raised to 25 degree celcius
pi1/pi2=t1v2/t2v1
500/105.3=283*v2/293*v2
500*293/105.3*283=v2/v1
v2/v1=146500/29799.9
v2/v1=4.916
v2/v1=5